Your line of mechanical pencils has monthly fixed costs of $28,000.

Each pencil has a variable cost of $1.85...

What is the break even point of these pencils sell of $2.70 each?
Answer #1

Fixed Cost= $28000

Variable cost = $1.85 per unit

Price= $2.70 per unit

Let assume that BEP is x units.

In breakeven point- total cost=total revenue

FC+VC*x= Price*x

28000+1.85*x= 2.70*x

2.70x-1.85x=28000

0.85x=28000

X= 32941.18 =32942units

Pl rate the answer